OpenStudy (anonymous):

write an equation for a line that passes through the point (-1,0) and is perpendicular to a line that has a slope of -2

OpenStudy (anonymous):

slope(line)*(-2)=-1, so slope(line)=1/2 y=mx+b 0=(1/2)(-1)+b b=1/2 y=(x/2) +1/2